"Write a quadratic function whose roots are 3 and -6."
We are doing the Zero Product Property. I need help because I don't know how to make a function based on roots.

Answer by Theo(13342)

you create a function from the roots as follows:

x = 3 is one of the roots.
x = 3 is an equation.
you want to convert the equation so that 0 is one one side of it.
you do that by subtracting 3 from both sides of the equation to get:
x - 3 = 0
(x-3) is one of your factors.

x = -6 is one of the roots.
x = -6 is an equation.
you want to convert the equation so that 0 is on one side of it.
you do that by adding 6 to both sides of the equation to get:
x + 6 = 0
(x+6) is the other one of your factors.

you multiply the factors together to get the function.
(x-3) * (x+6) = x^2 + 6x - 3x - 18
your function is x^2 + 3x - 18.